OpenStudy (anonymous):

The sides of an isosceles triangle are 5, 5, and 7. Find the measure of the vertex angle to the nearest degree.

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@luckycoins888

OpenStudy (perl):

you can law of cosines

OpenStudy (perl):

Let me get the equation started 5^2 + 5^2 - 2*5*5cos(theta) = 7^2

OpenStudy (anonymous):

49?

OpenStudy (perl):

not quite, did you solve that?

OpenStudy (perl):

5^2 + 5^2 - 2*5*5cos(theta) = 7^2 25 + 25 - 50 cos(theta) = 49 50 - 50*cos(theta) = 49 -50cos(theta) = 49 - 50 -50*cos(theta) = -1 cos(theta) = 1/50

OpenStudy (anonymous):

thnx